Laurel Enterprises purchased equipment for $72,000 on January 1, 2017. The equipment is expected to have a five-year life and a residual value of $6,000. Using this information & the double-declining balance method, depreciation expense for 2017 and the net book value at December 31st, 2017 is:

a. $26,400 & $36,600

b. $26,400 & $45,600

c. $28,800 & $37,200

d. $28,800 & $43,200

PLEASE SHOW WORK
Answer #1

Annual depreciation rate as per straight line =100/5=20%/year

Hence depreciation as per double decline=2*Annual depreciation rate as per straight line *Beginning value of each period

Year Beginning value Depreciation Ending value
2017 72000 (2*20%*72000)=$28800 (72000-28800)=$43200

Hence the correct option is D.